# Improving Low-Resource Sequence Labeling with Knowledge Fusion and Contextual Label Explanations

### v2

@@ 4 Experiment
Dataset / Model Weibo Youku Taobao Resume
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003) 56.57 60.91 66.52 68.02 70.57 74.92 68.78 71.88 74.74 90.19 92.35 93.43
FLAT (Li et al., 2020) 57.75 59.47 65.72 72.31 76.01 78.73 69.84 71.72 76.21 91.35 93.04 93.61
MECT (Wu et al., 2021) 58.55 60.77 66.13 72.82 75.85 79.16 70.54 73.87 76.48 91.52 93.63 93.90
LEBERT (Liu et al., 2021) 61.23 64.03 67.63 72.39 75.00 77.88 71.12 74.46 77.44 93.08 94.16 95.05
PCBERT (Lai et al., 2022) 70.73 70.78 72.81 77.67 81.96 83.66 73.32 75.41 79.21 93.63 94.31 95.18
BiaffineNER (Yu et al., 2020) 58.74 66.41 69.70 77.28 80.21 81.68 74.62 76.98 79.46 93.30 94.61 95.49
[…]
CNN Nested NER (Yan et al., 2023) 64.81 67.75 69.96 79.28 81.58 83.94 75.39 77.86 80.06 93.10 94.54 95.39
DiFiNet (Cai et al., 2024) 67.35 69.02 72.19 79.81 81.32 83.29 75.40 77.05 79.61 93.81 94.75 95.93
KnowFREE + Fusion (GLM3-6B) 66.76 71.59 72.99 79.30 82.13 84.50 76.31 78.55 80.53 94.03 95.04 96.14
@@
Dataset / Model CMeEE-v2 PKU MSR UD
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003) - - - 93.49 94.31 95.02 90.60 91.73 92.93 87.11 89.87 91.98
FLAT (Li et al., 2020) - - - 78.28 80.10 80.22 77.72 78.27 78.44 77.76 78.39 80.11
MECT (Wu et al., 2021) - - - 87.28 87.54 87.58 87.48 87.53 87.71 87.12 87.30 87.61
LEBERT (Liu et al., 2021) - - - 93.42 94.09 94.97 90.64 91.95 93.27 88.93 91.85 93.43
PCBERT (Lai et al., 2022) - - - 93.47 94.07 94.54 90.01 92.18 93.10 89.81 91.70 93.67
BiaffineNER (Yu et al., 2020) 61.26 65.56 68.40 93.20 94.39 94.94 91.60 92.63 93.64 88.84 90.74 92.68
W2NER (Li et al., 2022) 61.10 65.67 68.72 93.90 94.64 95.41 91.61 92.76 93.82 90.12 92.91 94.86
CNN Nested NER (Yan et al., 2023) 62.32 66.43 69.06 94.00 94.59 95.47 91.72 92.86 93.67 90.21 92.48 94.70
DiFiNet (Cai et al., 2024) 63.55 66.26 67.28 93.81 94.76 95.26 91.46 92.45 93.24 90.94 92.82 94.62
KnowFREE + Fusion (GLM3-6B) 63.64 67.47 69.52 94.07 94.94 95.51 91.73 92.91 93.92 90.99 92.71 95.00
@@
[…]
KnowFREE + Fusion (Qwen-14B) 63.23 67.25 69.19 94.36 94.94 95.51 91.76 92.93 93.89 90.53 93.03 94.97
[…]
Table 1: The overall results in few-shot sequence labeling tasks, where the value highlighted in bold is the best value, and the value underline is the second-best value.
@@ 4.1 Experiment Setup
Training: To comprehensively evaluate the effectiveness of our data augmentation strategy on different LLMs, and to fairly compare previous related methods. We use GLM3-6B (GLM et al., 2024), GLM4-9B (GLM et al., 2024) and Qwen2.5-14B (Yang et al., 2024b, a) as LLMs for label extension annotation and enriched explanation synthesis. And we choose BERT (Devlin et al., 2019) as the backbone model of KnowFREE encoder. In the pipeline of label extension annotation, the pre-trained embedding model of $\mathcal{M}$ is set as “text2vec” (Ming, 2022), the Top-$p$ value in Equation 7 is set as five, and the threshold $\epsilon$ is set as 1.5. In the training stage, the hidden size $D$, $D^{\prime}$, and $\tilde{D}$ are set as 768, 200, and 200, respectively. The activation function of $\sigma$ and $\sigma^{\ast}$ are defined as Leaky ReLU and GeLU, respectively. In the local multi-head attention module, the window size $\omega$ is set as three and the number of attention heads $K$ is set as ten. The detail settings of $\alpha$, $\beta$ and other parameter are presented in Appendix G.
Evaluation: To evaluate the performance of our method in low-resource sequence labeling tasks, we conducted experiments on multiple datasets. These include four flat NER datasets: Weibo (Peng and Dredze, 2015), Youku (Jie et al., 2019), Taobao (Jie et al., 2019), and Resume (Zhang and Yang, 2018b); a nested NER dataset, CMeEE-v2 (Zhang et al., 2022); two word segmentation datasets, PKU (Emerson, 2005) and MSR (Emerson, 2005); and a POS tagging dataset, UD (Nivre et al., 2016). Real-world datasets often exhibit imbalanced category distributions, making them ideal for evaluating a model’s generalization ability in authentic low-resource scenarios. In our main experiments, we randomly sampled subsets of varying sizes from the datasets to simulate sequence labeling under low-resource conditions. The training set sample sizes $\mathcal{K}$ were set to 250, 500, and 1000, respectively. The detail statistics of all datasets are presented in Appendix F. We control randomness by ensuring that subsequent samples include previously sampled samples.
Baselines: To provide a fair comparison of the effectiveness of our method, we selected several mainstream approaches for Chinese sequence labeling tasks, including BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003), FLAT (Li et al., 2020), MECT (Wu et al., 2021), and LEBERT (Liu et al., 2021). Additionally, we included PCBERT (Lai et al., 2022) that also leverage extension knowledge enhancement, as well as methods designed for nested entity recognition, including BiaffineNER (Yu et al., 2020), W2NER (Li et al., 2022), CNN Nested NER (Yan et al., 2023), and DiFiNet (Cai et al., 2024). In our analysis, we further compared the performance of vanilla LLMs and LoRA fine-tuned models on sequence labeling tasks. Additionally, we evaluated the performance of various LLM-based data augmentation strategies.
@@ 4.2 Main Results
As shown in Table 1, our method consistently outperforms others on average. On the Weibo dataset, our best result surpasses PCBERT by 6.4%. Under the $\mathcal{K}=250$ setting, our approach outperforms the strongest baseline by an average margin of 1.95% across all datasets. Furthermore, incorporating the enriched explanation synthesis pipeline (+ Fusion + Synthetic) improves performance over using only label extension annotation (+ Fusion) when $\mathcal{K}$=250 across all datasets. This effect is particularly evident on the Weibo dataset, where our best GLM3-6B-based “+ Fusion + Synthetic” ($\mathcal{K}$=1000) surpasses the performance of “+ Fusion” ($\mathcal{K}$=1000) by 1.79%. However, as $\mathcal{K}$ increases, the performance gains from “+ Fusion + Synthetic” in most datasets gradually decline, and fall slightly below “+ Fusion”. This suggests that enriched data synthesis is more effective in resource-scarce scenarios. Meanwhile, under the “+ Fusion ($\mathcal{K}$=1000)” setting, our best results achieves an average improvement of 0.52% over the strongest baseline across all NER datasets, further validating the effectiveness of the label extension annotation strategy. Finally, we observe that different LLM backbones yield varying performance across datasets. While larger LLMs contribute more significantly to the improvement of “+ Fusion” in low-resource scenarios, they do not exhibit a clear advantage in data synthesis. This may be due to variations in how different LLMs align with prompt instructions. Since our prompts were designed and validated on GLM3-6B, their effectiveness may differ when applied to other LLMs.
@@ 4.3 Analysis
Ablation Studies: To assess the impact of each module and strategy in our method, we conducted ablation experiments by removing individual components and evaluating their effects on model performance. The results are summarized in Table 2. “w/o L.E.A.” refers to the model’s performance without the label extension annotation, relying solely on the vanilla KnowFREE model. This configuration results in a significant performance drop compared to the default method. Nevertheless, it still outperforms W2NER and CNN Nested NER on datasets such as Weibo, Youku, CMeEE-v2, and UD, demonstrating the effectiveness of the local multi-head attention module in capturing and enhancing neighborhood feature interactions. “w/o S.L.” evaluates the impact of removing the synonymous label merging strategy, which results in severe performance degradation. This highlights that even with the inclusion of extension label features, the absence of synonymous label merging leads to confusion when equivalent labels are treated as distinct categories, impairing the model’s judgment. “w/o entity” and “w/o pos,” evaluate the removal of extension entities and POS tagging features, respectively. On NER datasets, the absence of extension entity features causes a more significant performance drop compared to POS tagging, highlighting the critical role of entity-related features in improving the model’s ability to detect entities. The impact of “w/o pos” varies across datasets. On MSR and UD, removing extension POS tagging features leads to improved performance. This may be due to noise introduced by the extension features. Since these features are highly correlated with the target labels, any noise in the extension features could disproportionately affect the model’s performance on the target task.
@@
Table 2: Results of F1 scores in ablation studies, all results are trained on datasets with $\mathcal{K}$=1000. The backbone LLM is GLM3-6B.
@@
Impact of Enriched Explanation Synthesis in K-Shot Sampling: To evaluate the impact of enriched explanation synthesis on model performance in low-resource scenarios, we conducted experiments following the n-way k-shot paradigm. Using a greedy sampling strategy, we repeatedly sampled data from flat NER datasets, ensuring that each target label in the k-shot setting was represented by at least k instances. The sampled data was then augmented with GLM3-6B. We compared the model’s performance with and without enriched explanation synthesis, as shown in Figure 4. Here, “w Synthetic” indicates the performance with enriched explanation synthesis, while “w/o Synthetic” reflects the performance without it. The results demonstrate that “w Synthetic” achieves a substantial performance boost over “w/o Synthetic” from the outset. Notably, when k is less than 15, enriched explanation synthesis consistently delivers rapid performance improvements across all datasets. As k increases, the performance gap narrows, but “w Synthetic” continues to outperform “w/o Synthetic” across all settings. These findings highlight that in resource-scarce scenarios, synthesizing enriched data is more effective than directly injecting features into raw samples. These findings highlight the critical role of enriched explanation synthesis in enhancing model performance, particularly when labeled data is limited.
Besides, we conduct extra ablation studies to analyze the effect of the number of heads in local attention in Appendix C. Additionally, we evaluate sequence labeling performance with LLMs in Appendix A and compare various data synthesis strategies in Appendix D. Finally, we visualize the logit scores of extension labels in Appendix E.

### v4

@@ 4 Experiment
Weibo Youku Taobao Resume CMeEE-v2 PKU MSR UD
Model 250 500 1000 250 500 1000 250 500 1000 250 500 1000 250 500 1000 250 500 1000 250 500 1000 250 500 1000
BERT-CRF (Devlin et al., 2019) 56.57 60.91 66.52 68.02 70.57 74.92 68.78 71.88 74.74 90.19 92.35 93.43 - - - 93.49 94.31 95.02 90.60 91.73 92.93 87.11 89.87 91.98
FLAT (Li et al., 2020) 57.75 59.47 65.72 72.31 76.01 78.73 69.84 71.72 76.21 91.35 93.04 93.61 - - - 78.28 80.10 80.22 77.72 78.27 78.44 77.76 78.39 80.11
MECT (Wu et al., 2021) 58.55 60.77 66.13 72.82 75.85 79.16 70.54 73.87 76.48 91.52 93.63 93.90 - - - 87.28 87.54 87.58 87.48 87.53 87.71 87.12 87.30 87.61
LEBERT (Liu et al., 2021) 61.23 64.03 67.63 72.39 75.00 77.88 71.12 74.46 77.44 93.08 94.16 95.05 - - - 93.42 94.09 94.97 90.64 91.95 93.27 88.93 91.85 93.43
PCBERT (Lai et al., 2022) 70.73 70.78 72.81 77.67 81.96 83.66 73.32 75.41 79.21 93.63 94.31 95.18 - - - 93.47 94.07 94.54 90.01 92.18 93.10 89.81 91.70 93.67
BiaffineNER (Yu et al., 2020) 58.74 66.41 69.70 77.28 80.21 81.68 74.62 76.98 79.46 93.30 94.61 95.49 61.26 65.56 68.40 93.20 94.39 94.94 91.60 92.63 93.64 88.84 90.74 92.68
W2NER (Li et al., 2022) 54.57 63.21 71.09 79.20 81.40 83.28 74.68 76.80 79.71 94.39 95.82 96.35 61.10 65.67 68.72 93.90 94.64 95.41 91.61 92.76 93.82 90.12 92.91 94.86
CNN Nested NER (Yan et al., 2023) 64.81 67.75 69.96 79.28 81.58 83.94 75.39 77.86 80.06 93.10 94.54 95.39 62.32 66.43 69.06 94.00 94.59 95.47 91.72 92.86 93.67 90.21 92.48 94.70
DiFiNet (Cai et al., 2024) 67.35 69.02 72.19 79.81 81.32 83.29 75.40 77.05 79.61 93.81 94.75 95.93 63.55 66.26 67.28 93.81 94.76 95.26 91.46 92.45 93.24 90.94 92.82 94.62
KnowFREE-F (ChatGLM3-6B) 66.76 71.59 72.99 79.30 82.13 84.50 76.31 78.55 80.53 94.03 95.04 96.14 63.64 67.47 69.52 94.07 94.94 95.51 91.73 92.91 93.92 90.99 92.71 95.00
KnowFREE-F (GLM4-9B-Chat) 66.40 73.08 72.59 79.40 82.16 84.37 76.02 78.21 80.48 94.05 95.43 96.25 63.98 67.41 69.38 94.57 95.01 95.49 91.83 92.73 93.91 90.58 92.72 94.98
KnowFREE-F (Qwen-14B) 68.07 73.39 73.41 80.30 82.10 84.20 76.38 78.00 80.50 94.21 95.32 96.40 63.23 67.25 69.19 94.36 94.94 95.51 91.76 92.93 93.89 90.53 93.03 94.97
@@
[…]
KnowFREE-FS (Llama3.1-70B-Instruct) 74.18 76.91 76.68 81.69 83.09 84.36 76.48 77.91 79.69 94.86 95.46 96.21 68.12 68.45 68.85 94.62 95.02 95.47 92.16 93.02 93.66 92.68 93.76 94.23
@@
Table 1: The overall results on many-shot sequence labeling tasks. KnowFREE-F denotes the variant using only the label extension annotation pipeline, while KnowFREE-FS incorporates the enriched explanation synthesis pipeline. The bold values indicate the best performance, and the underlined values represent the second-best.
@@ 4.1 Experiment Setup
Training: To comprehensively evaluate the effectiveness of our data augmentation strategy on different LLMs, and to fairly compare previous related methods. We use ChatGLM3-6B, GLM4-9B-Chat, Qwen2.5-14B-Instruct, Llama3.1-70B-Instruct, GPT-4o and Deepseek-V3 (GLM et al., 2024; Yang et al., 2024b, a; Dubey et al., 2024; OpenAI, 2023; DeepSeek-AI et al., 2025) as LLMs for label extension annotation and enriched explanation synthesis. We choose BERT (Devlin et al., 2019) as the backbone encoder of the KnowFREE. More detail settings are presented in Appendix H.
Evaluation: To assess the performance of our method in low-resource scenarios, we conducted experiments on a variety of datasets. These include Chinese flat NER datasets (Weibo (Peng and Dredze, 2015), Youku (Jie et al., 2019), Taobao (Jie et al., 2019), and Resume (Zhang and Yang, 2018b)); English flat NER datasets (CoNLL’03 (Sang and De Meulder, 2003) and MIT-Movie (Liu et al., 2013)); a Chinese nested NER dataset (CMeEE-v2 (Zhang et al., 2022)); word segmentation datasets (PKU and MSR (Emerson, 2005)); and a POS tagging dataset (UD (Nivre et al., 2016)). To evaluate our method under data scarcity and explore the limits of performance gains from sample synthesis, we conducted both many-shot and few-shot experiments. In the many-shot setting, we simulated low-resource conditions by randomly sampling subsets of 250, 500, and 1000 training instances. In the few-shot setting, we adopted the standard “n-way k-shot” paradigm, using a greedy sampling strategy to ensure each target label appeared at least $k$ times. To ensure consistency, each larger subset included all samples from the smaller ones. Dataset statistics are provided in Appendix G. We then discuss the effectiveness of each module in the analysis section. Additionally, we report the results on the full datasets in Appendix B, conduct further ablation studies on the number of heads in local attention in Appendix D, and provide visualizations of the logit scores for the extension labels in Appendix F.
Baselines: To ensure a fair comparison, we evaluate our method against both model-centric and data-centric baselines. On the model-centric side, we compare with general baselines such as BERT-CRF (Devlin et al., 2019) and BiaffineNER (Yu et al., 2020), as well as models specifically designed for Chinese sequence labeling, including FLAT (Li et al., 2020), MECT (Wu et al., 2021), and LEBERT (Liu et al., 2021). We also include comparisons with state-of-the-art nested NER methods such as W2NER (Li et al., 2022), CNN Nested NER (Yan et al., 2023), and DiFiNet (Cai et al., 2024). From the data-centric perspective, we compare with the transfer learning approach PCBERT (Lai et al., 2022), and LLM-enhanced methods including LLM-DA (Ye et al., 2024), ProgGen (Heng et al., 2024), and MELM (Zhou et al., 2022). In addition, Appendix A provides results comparing vanilla LLMs and LoRA fine-tuned models on sequence labeling tasks.
@@ 4.2 Main Results
Many-shot Results. As shown in Table 1, our method consistently achieves higher average performance across all datasets. We observe that larger backbone LLMs generally bring greater performance improvements to our approach. Although the scaling law is not strictly linear, even the smallest model, ChatGLM3-6B, delivers strong results. Under the 250-sample setting, our method surpasses the strongest baseline by an average of 1.95%, especially with a 4.05% gain on the Weibo dataset. In low-resource settings, KnowFREE-FS outperforms KnowFREE-F. However, as the number of training samples increases, especially beyond 500, the performance of KnowFREE-FS becomes comparable to or slightly lower than that of KnowFREE-F. This indicates that enriched data synthesis is more effective when training data is limited. When more data is available, the noise introduced by synthetic samples may outweigh the benefits. Further analysis of noise effects is provided in Appendix E. Even when data synthesis becomes less effective in higher-resource scenarios, KnowFREE-F maintains strong performance. With 1000 training samples, it still outperforms the strongest baseline by an average of 0.95% across all NER datasets, demonstrating the robustness and effectiveness of the label extension annotation strategy.
Size Ours LLM-DA ProgGen MELM C.N.-NER DiFiNet
Weibo
kk=5 35.58 (2.33) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
kk=10 48.53 (1.57) 0.95 (0.23) 0.48 (0.19) 0.03 (0.04) 0.95 (0.21) 0.98 (0.18)
kk=15 60.61 (2.58) 17.28 (1.93) 12.36 (1.22) 2.98 (0.58) 20.68 (1.88) 27.43 (2.30)
kk=20 68.62 (2.04) 39.49 (2.41) 34.16 (2.09) 16.71 (1.67) 31.10 (2.01) 32.71 (2.31)
Youku
kk=5 38.76 (2.30) 12.03 (2.24) 13.76 (1.32) 9.98 (2.01) 24.70 (2.55) 23.83 (2.52)
kk=10 68.95 (3.85) 33.10 (2.34) 33.52 (2.01) 16.17 (1.30) 46.39 (2.48) 46.62 (2.43)
kk=15 71.76 (3.33) 59.72 (3.94) 56.55 (1.56) 50.18 (1.67) 60.41 (1.88) 60.61 (1.91)
kk=20 72.83 (2.29) 64.58 (2.53) 61.80 (1.60) 52.38 (1.59) 67.38 (1.92) 68.38 (1.58)
Taobao
kk=5 62.96 (2.34) 13.95 (2.32) 22.74 (2.62) 8.97 (0.88) 22.40 (1.61) 23.82 (1.56)
kk=10 64.64 (3.58) 54.05 (2.92) 50.75 (1.66) 32.41 (1.37) 53.33 (2.01) 53.75 (2.03)
kk=15 69.26 (2.65) 59.69 (2.58) 59.77 (1.51) 55.32 (1.57) 60.45 (1.83) 61.01 (1.54)
kk=20 68.93 (2.36) 61.79 (1.51) 61.77 (1.63) 43.13 (1.62) 63.36 (1.86) 64.08 (1.53)
Resume
kk=5 65.28 (0.97) 30.44 (0.66) 27.63 (0.34) 20.67 (1.26) 25.50 (1.22) 29.97 (1.17)
kk=10 78.89 (0.53) 45.40 (0.58) 50.39 (0.82) 42.34 (1.28) 49.78 (1.53) 50.19 (1.49)
kk=15 85.43 (1.17) 58.77 (1.15) 60.32 (1.26) 53.18 (1.18) 56.04 (1.31) 56.92 (1.16)
kk=20 85.56 (1.11) 75.13 (1.36) 82.96 (1.28) 69.15 (1.21) 67.51 (1.17) 68.83 (1.19)
CMeEE-v2
kk=5 49.68 (1.89) 35.03 (1.65) 39.18 (1.52) 12.78 (1.01) 6.49 (0.56) 5.62 (0.47)
kk=10 60.46 (1.67) 48.91 (1.61) 44.80 (1.59) 32.76 (1.65) 47.40 (1.56) 47.25 (1.50)
kk=15 62.46 (1.51) 48.97 (1.54) 49.32 (1.61) 38.79 (1.53) 48.90 (1.65) 48.72 (1.59)
kk=20 63.83 (1.68) 57.18 (1.63) 57.76 (1.51) 50.12 (1.67) 56.38 (1.58) 56.22 (1.55)
CoNLL’03
kk=5 64.18 (1.62) 57.84 (1.93) 57.11 (2.56) 30.00 (2.13) 27.75 (1.61) 26.86 (1.36)
kk=10 75.83 (1.52) 69.07 (2.44) 69.10 (2.39) 63.48 (2.18) 62.93 (1.94) 59.17 (1.88)
kk=15 78.68 (1.39) 78.18 (2.22) 78.32 (2.19) 76.16 (2.03) 75.93 (1.86) 75.85 (1.89)
kk=20 83.24 (1.21) 81.94 (2.17) 82.09 (1.55) 79.41 (2.21) 77.92 (1.58) 77.74 (1.81)
MIT-Movie
kk=5 57.34 (1.88) 53.97 (2.07) 52.82 (2.44) 38.49 (2.23) 36.81 (1.26) 37.62 (1.33)
kk=10 64.08 (1.66) 63.03 (2.35) 63.41 (2.14) 50.56 (2.29) 49.08 (1.60) 49.43 (1.58)
kk=15 67.03 (1.68) 65.77 (1.46) 65.93 (1.17) 58.32 (1.87) 58.03 (1.69) 58.54 (1.51)
kk=20 69.28 (1.63) 69.12 (1.08) 69.19 (1.59) 62.02 (1.91) 60.81 (1.60) 61.07 (1.64)
Table 2: Results of few-shot sequence labeling tasks. Our default method is KnowFREE-FS (ChatGLM3-6B). C.N.-NER refers to the abbreviation of CNN Nested NER. Values in parentheses indicate standard deviation. bold numbers highlight the best performance.
Few-shot Results. To assess the effectiveness of our method in few-shot settings, we compared it with state-of-the-art nested NER models and several LLM-based data augmentation strategies, including LLM-DA, ProgGen, and MELM, on both Chinese and English NER datasets. For fair comparison, all synthesized samples were annotated using the KnowFREE model trained solely on the original data, and the resulting data were used to retrain KnowFREE. As shown in Table 2, our method consistently outperforms the baselines under few-shot settings. On the Weibo dataset with kk=5, while other methods yield zero performance, our approach achieves the performance of 35.58%. Moreover, for $k\leq 15$, LLM-based augmentation strategies often perform worse than CNN Nested NER and DiFiNet, indicating limited domain adaptability and the adverse effects of noise introduced by data synthesis. The performance gains are more pronounced on Chinese datasets compared to English ones, demonstrating the method’s robustness across languages and its particular strength in character-dense languages. Further analysis of performance trends of LLM-based methods under varying data sizes is provided in Appendix E.
@@ 4.3 Analysis
Table 3: Results of F1 scores in ablation studies, all results are trained on datasets with 1000. The backbone LLM is ChatGLM3-6B.
Ablation Studies: To evaluate the contribution of each component in our approach, we conducted ablation studies by selectively removing modules and analyzing their impact on model performance, as shown in Table 3. “w/o L.E.A.” removes the Label Extension Annotation module and uses the vanilla KnowFree model. Although this leads to a performance drop, it still outperforms nested NER baselines across several datasets. “w/o local attn & L.E.A.” disables both the local attention and L.E.A. modules, resulting in a significant average performance drop of 1.56%, highlighting their combined effectiveness. In “w/o local attn w CNN,” the local attention module is replaced by the masked CNN module from CNN Nested NER. While this improves performance over CNN Nested NER, it underperforms compared to our attention-based model, confirming the advantage of local multi-head attention for capturing neighborhood interactions. “w/o S.L.” removes the synonymous label merging strategy and causes a 0.42% drop in performance, indicating that failing to unify semantically equivalent labels introduces confusion and weakens model predictions. In “w/o entity” and “w/o pos,” we exclude extension entity features and POS features, respectively. Removing entity features leads to a larger performance drop, showing their stronger impact on entity recognition. Interestingly, removing POS features improves results on MSR and UD, possibly due to noise introduced by imperfect or overly correlated POS tags.
@@
Figure 5: t-SNE visualization of the training, test and enriched explanation samples under different sampling sizes. The synthetic enriched explanation samples are generated by ChatGLM3-6B, and they are represented by the “Synthetic” in the legend.
Impact of Enriched Explanation Synthesis in K-Shot Sampling: To further evaluate the impact of enriched explanation synthesis on model performance in low-resource scenarios, we conducted experiments following the “n-way k-shot” paradigm. The sampled data was then augmented with ChatGLM3-6B. We compared the model’s performance with and without enriched explanation synthesis, as shown in Figure 4. Here, “w Synthetic” indicates the performance with enriched explanation synthesis, while “w/o Synthetic” reflects the performance without it. The results demonstrate that “w Synthetic” achieves a substantial performance boost over “w/o Synthetic” from the outset. Notably, when k is less than 15, enriched explanation synthesis consistently delivers rapid performance improvements across all datasets. As k increases, the performance gap narrows, but “w Synthetic” continues to outperform “w/o Synthetic” across all settings. These findings highlight that in resource-scarce scenarios, synthesizing enriched data is more effective than directly injecting features into raw samples. These findings highlight the critical role of enriched explanation synthesis in enhancing model performance, particularly when labeled data is limited.
Visual Analysis of Enriched Explanation Synthesis: We use the text2vec embedding model (Ming, 2022) to generate sentence embeddings for the original training and test samples, as well as for the enriched explanation samples from the Weibo, Youku, Taobao, Resume, and CMeEE-v2 datasets. These embeddings are projected into two dimensions using t-SNE, with results shown in Figure 5. At a sample size of 250, the training data in datasets such as Weibo, Youku, Taobao, and Resume provides sparse semantic coverage, leaving portions of the test set insufficiently represented. This limitation is observed across all datasets. The synthesized samples help bridge these gaps in semantic space, which explains the substantial performance gains in low-resource scenarios. As the number of training samples increases, the coverage of the semantic space becomes more comprehensive for most datasets. However, at a sample size of 1000, semantic discrepancies appear between training and synthesized samples in datasets such as Youku, Taobao, and CMeEE-v2, which may introduce noise and hinder model performance. In the Weibo dataset, certain regions of the test semantic space remain underrepresented even at this larger sample size. This observation explains why models trained with synthesized samples continue to outperform those trained solely on original data at this sample size.
These findings underscore the effectiveness of enriched explanation synthesis in enhancing model performance under low-resource conditions. However, they also reveal that as the sample size grows, the potential drawbacks of synthesized data, such as semantic noise, become increasingly evident.